# GraphPFN: A Prior-Data Fitted Graph Foundation Model

### v1

@@ Abstract
Foundation models pretrained on large-scale datasets have transformed such fields as natural language processing and computer vision, but their application to graph data remains limited. Recently emerged graph foundation models, such as G2T-FM, utilize tabular foundation models for graph tasks and were shown to significantly outperform prior attempts to create GFMs. However, these models primarily rely on hand-crafted graph features, limiting their ability to learn complex graph-specific patterns. In this work, we propose GraphPFN: a prior-data fitted network for node-level prediction. First, we design a prior distribution of synthetic attributed graphs. For graph structure generation, we use a novel combination of multiple stochastic block models and a preferential attachment process. We then apply graph-aware structured causal models to generate node attributes and targets. This procedure allows us to efficiently generate a wide range of realistic graph datasets. Then, we augment the tabular foundation model LimiX with attention-based graph neighborhood aggregation layers and train it on synthetic graphs sampled from our prior, allowing the model to capture graph structural dependencies not present in tabular data. On diverse real-world graph datasets with up to 50,00050{,}000 nodes, GraphPFN shows strong in-context learning performance and achieves state-of-the-art results after finetuning, outperforming both G2T-FM and task-specific GNNs trained from scratch on most datasets. More broadly, our work demonstrates that pretraining on synthetic graphs from a well-designed prior distribution is an effective strategy for building graph foundation models. [^1] Our code is available at https://github.com/yandex-research/graphpfn.
@@ 1 Introduction
In recent years, foundation models have significantly advanced the state of the art in domains such as natural language processing and computer vision. These models can learn from large unannotated datasets and generalize across a wide range of downstream tasks. Notable examples include large language models like BERT (Devlin et al., 2019) and GPT-3 (Brown et al., 2020) in natural language processing and the Vision Transformer (Dosovitskiy et al., 2020) and CLIP (Radford et al., 2021) in computer vision. These models often use self-supervised or unsupervised pretraining to learn rich, transferable representations. This approach has changed how models are built and used, removing the need for task-specific models and reducing dependence on large labeled datasets for every single task. Inspired by these successes, there is growing interest now in extending the foundation models methodology to other modalities, including graphs.
@@
However, developing graph foundation models (GFMs) is much more challenging. Unlike text and images, graph data does not constitute a single domain. Instead, graphs are used to represent data from different domains, e.g., social networks (both virtual and real-world), information networks, transportation networks, co-purchasing networks, various physical, biological, or engineering systems, or even networks of abstract concepts. As a consequence, both the structure of a graph and its attributes (features and labels) may vary significantly across graph datasets and tasks. This makes it difficult to design truly generalizable graph foundation models.
@@
However, it has recently been noted by Eremeev et al. (2025) that the problem of feature and target space heterogeneity, i.e., the use of different feature and target spaces for different graph datasets, is not unique to graph machine learning, but is also faced by tabular machine learning which deals with similarly diverse datasets from very different domains. To address this problem, recent tabular foundation models (TFMs) employ the framework of prior-data fitted networks (PFNs, Müller et al., 2021; Hollmann et al., 2023). PFNs are pretrained on diverse synthetic datasets to approximate Bayesian inference and can make predictions in an in-context learning setting. Recent works, such as TabPFNv2 (Hollmann et al., 2025) or LimiX (Zhang et al., 2025), show that this approach enables the creation of strong tabular foundation models.
@@
Motivated by the success of tabular foundation models, recent works G2T-FM (Eremeev et al., 2025) and TabGFM (Hayler et al., 2025) have utilized foundation models for tabular data to create graph foundation models. For this, they augment node features with graph-based information such as neighborhood-aggregated features or Laplacian positional encodings. This allows for transforming a graph node-level prediction problem into a tabular prediction problem and applying an existing tabular foundation model to this task. The resulting approach shows strong performance, but such models still depend heavily on hand-crafted features and, as a result, are limited in their ability to capture complex graph patterns.
To address these limitations, we propose GraphPFN: a Graph Prior-data Fitted Network. Unlike previous approaches, GraphPFN does not rely on hand-crafted features. Instead, it includes a trainable message-passing mechanism standard for graph neural networks (GNNs), which allows it to model more complex graph dependencies. Specifically, we initialize GraphPFN from the tabular foundation model LimiX (Zhang et al., 2025) and add an attention-based message-passing layer to each block of LimiX. This allows the model to learn graph-specific patterns while keeping its ability to handle diverse features and labels inherited from LimiX.
@@
We pretrain GraphPFN on synthetic datasets using the PFN framework. To create these datasets, we introduce a novel graph prior. For generating graph structures, we propose an approach that combines multiple stochastic block models and augments them with a preferential-attachment process. We then generate graph-structure-dependent node attributes for our graphs by augmenting tabular structured causal models (SCMs, Hollmann et al., 2023; Qu et al., 2025) typical for tabular PFNs with message-passing mechanisms at random SCM nodes. This method allows us to efficiently generate millions of realistic and diverse synthetic graph datasets.
@@
Our experiments show that on diverse real-world graph datasets with up to 50,00050{,}000 nodes, GraphPFN achieves strong in-context learning performance, competitive with the best current models — well-tuned traditional GNNs (Kipf & Welling, 2017; Hamilton et al., 2017; Veličković et al., 2018; Shi et al., 2021) with improved architectures (Platonov et al., 2023) and G2T-FM (Eremeev et al., 2025). Furthermore, after finetuning, GraphPFN outperforms all other approaches, setting a new state-of-the-art for the considered datasets.
@@
  • We introduce a novel graph prior for efficient generation of realistic synthetic attributed graphs.
  • We propose GraphPFN, which is, to the best of our knowledge, the first publicly available successful implementation of PFN framework for graph data.
  • We demonstrate that finetuned GraphPFN outperforms both strong traditional GNN baselines and existing graph foundation models.
@@ 2.2 Tabular foundation models
In their pioneering work TabPFN (Hollmann et al., 2023) and its successor TabPFNv2 (Hollmann et al., 2025), the authors proposed to utilize the framework of prior-data fitted networks to create tabular foundation models and showed that such models can achieve strong results, competitive with other approaches (Erickson et al., 2025). Nowadays, TFMs have become an active research area, with several new methods released recently (Zhang & Robinson, 2025; Zhang et al., 2025). Some methods focus on scalability (Qu et al., 2025) or faster inference (Mueller et al., 2025), while others emphasize training on real-world datasets rather than synthetic tasks (Ma et al., 2024). Together, these works broaden the design space of tabular foundation models by trading off data sources, computational efficiency, and scalability.
@@ 2.3 Graph foundation models
Similar to foundation models for tabular data, graph foundation models also face the challenge of handling datasets from diverse domains. A particularly difficult, yet crucial, aspect is managing the wide variety of node attributes (features and labels) present in different graphs. Early GFMs did not fully address this issue. They often relied on dimensionality reduction techniques such as principal component analysis or singular value decomposition (Xia & Huang, 2024; Zhao et al., 2024; Wang et al., 2025; Yu et al., 2025), or they restricted their focus to graphs where node attributes are all textual (Wang et al., 2024; He & Hooi, 2024; Liu et al., 2024).
More recent works, such as G2T-FM (Eremeev et al., 2025) and TabGFM (Hayler et al., 2025), have explored leveraging tabular foundation models to better address feature diversity in graph datasets. These approaches incorporate hand-crafted features, for example, neighborhood-aggregated features or Laplacian positional encodings, to effectively convert graph information into tabular features. Empirical results show that these methods achieve strong results, always significantly outperforming prior GFMs (Xia et al., 2024; Xia & Huang, 2024; Zhao et al., 2024; Finkelshtein et al., 2025; Zhao et al., 2025) and frequently outperforming well-tuned GNNs trained from scratch (Kipf & Welling, 2017; Hamilton et al., 2017; Veličković et al., 2018; Shi et al., 2021) with improved architectures (Platonov et al., 2023), supporting the utility of employing tabular foundation models as a basis for learning on graph data.
@@ 3.1 Architecture
Our model architecture extends the tabular foundation model LimiX (Zhang et al., 2025) by adding attention-based message-passing layers to each of its transformer blocks. This strategy is inspired by recent findings that tabular foundation models can already learn patterns relevant for various graph tasks (Eremeev et al., 2025; Hayler et al., 2025). By initializing our model with a pretrained tabular foundation model instead of training from scratch, we leverage these learned representations, which significantly reduces computational costs while still achieving strong results.
Below we first summarize the LimiX backbone to clarify how GraphPFN represents samples (nodes) and features, and how attention flows in the base model. We then describe how the graph adapters modify this flow to leverage the graph topology.
@@
LimiX (Zhang et al., 2025) is a transformer-style foundation model for tabular data that departs from the common design of representing each sample with a single fixed-length embedding. Instead, it uses a multi-token representation: for every sample, each feature contributes one token, [^2] In the current implementation, two features are grouped into one token, but we omit this detail for clarity. yielding a token grid with one axis for features and one for samples. This design naturally handles a variable number of heterogeneous features in different datasets without changing and re-training the model.
A LimiX transformer block contains three attention layers, each followed by a element-wise feed-forward network (FFN). Two layers are feature-level multi-head attention (MHA) modules that operate within a sample, allowing all feature tokens of the same sample to attend to each other. And the third is a sample-level MHA that operates within a feature across samples, allowing tokens corresponding to the same feature to exchange information across the dataset. The feature-level MHAs enable rich interactions among features within each sample, while the sample-level MHA supports in-context learning by transferring information across samples for the same feature.
@@
To inject graph structure information without disrupting the LimiX’s tokenization, we add a message-passing adapter that implements scaled dot product attention between neighboring nodes to the end of every LimiX transformer block (see Figure 1 for an illustration). Note that we use the scaled dot product attention that is identical to the original Transformer attention (Vaswani et al., 2017) except for being restricted to 1-hop graph neighborhoods. Intuitively, our message-passing adapter performs a second, graph-structure-aware round of sample-level attention: tokens may exchange information only along the observed edges. In contrast to the global sample-level attention of LimiX (which is masked by the PFN protocol), the graph adapter is masked by the adjacency and therefore routes information locally, from each node to its neighbors. Because we keep the per-feature token representation intact, the adapter runs over the sample axis for each feature token independently, using the same graph mask across features.
@@
Each adapter is implemented as a sparse, multi-head attention module with the adjacency as its mask (two nodes attend if and only if an edge connects them). Similar to other attention modules, it is followed by a feed-forward network independently applied to each token. Both FFN and message-passing layer are wrapped with residual connections (He et al., 2016) and layer normalization (Ba et al., 2016), mirroring the structure of the LimiX blocks for stable optimization.
@@ 3.2 Pretraining
GraphPFN was pretrained following the PFN framework (Müller et al., 2021; Hollmann et al., 2023; 2025), using continuous pretraining from the LimiX checkpoint. Specifically, we generated 4,000,0004{,}000{,}000 synthetic datasets according to the prior described below and used them for pretraining. The pretraining process was conducted for 500,000500{,}000 steps on 8 NVIDIA A100 80G GPUs, with each GPU processing one synthetic dataset per step. The total pretraining time was approximately 6 days.
To monitor the progress of pretraining, we evaluated the in-context learning performance of GraphPFN every 100 steps on several datasets from the GraphLand benchmark (Bazhenov et al., 2025) (see Section 5.1 for the evaluation procedure). We emphasize that these evaluations were used only for a post-hoc analysis of training progress: test set performance was not used for early stopping or any other form of model selection. The evaluation results can be found in Figure 2.
@@
We optimize a joint objective that combines the PFN supervised loss with the masked graph modeling (MGM) loss (Li et al., 2023). For the supervised PFN term, for each dataset we sample a random set of context nodes, compute predictions for all other nodes, and minimize the cross-entropy loss on them.
To encourage more graph-aware representations, inspired by the success of GNN self-supervised pretraining, we add the masked graph modeling term from Li et al. (2023). Specifically, we randomly sample a fraction $p=0.1$ of edges as positive samples, remove these edges from the input graph, and uniformly sample an equal number of unconnected node pairs as negative examples. We then train the model with the cross-entropy loss to distinguish positive and negative edges. For this purpose, we add an additional MLP head on top of the target embeddings from the last layer. The total loss is the sum of the supervised loss and the MGM loss, with a coefficient of $0.1$ applied to the MGM term.
Details
To ensure training stability and retain the strong feature modeling capabilities of LimiX, we froze all model layers except for the graph adapters, which were the only components updated during pretraining. Optimization was performed using the Adam optimizer (Kingma & Ba, 2014) with a constant learning rate of $\gamma=3\cdot 10^{-4}$ and a linear warmup schedule over the first 10,00010{,}000 steps.
Figure 2: In-context learning performance of intermediate checkpoints of GraphPFN. The $x$-axis represents the number of steps, and the $y$-axis represents the metric on the test set.

### v2

@@ Abstract
Graph foundation models face several fundamental challenges including transferability across datasets and data scarcity, which calls into question the very feasibility of graph foundation models. However, despite similar challenges, the tabular domain has recently witnessed the emergence of the first successful foundation models such as TabPFNv2 and LimiX. Many of these models are based on the prior-data fitted networks (PFN) framework, in which models are pretrained on carefully designed synthetic datasets to make predictions in an in-context learning setting. Recently, G2T-FM has made the first step towards adopting PFNs for graphs, yet it is limited to hand-crafted features and was never pretrained on graph data. In this work, we make the next step by proposing GraphPFN, a PFN-based model designed and pretrained specifically for graph node-level tasks. Following the PFN framework, we first design a prior distribution of synthetic attributed graphs by using a novel combination of multi-level stochastic block models and a preferential attachment process for structure generation and graph-aware structured causal models for attribute generation. Then, we augment the tabular foundation model LimiX with attention-based graph neighborhood aggregation layers and train it on synthetic graphs sampled from our prior. On diverse real-world graph datasets with node-level tasks, GraphPFN shows strong in-context learning performance and achieves state-of-the-art results after finetuning, outperforming both G2T-FM and task-specific GNNs trained from scratch on most datasets. More broadly, GraphPFN shows the potential of PFN-based models for building graph foundation models. Our code is available at https://github.com/yandex-research/graphpfn.
@@ 1 Introduction
In recent years, foundation models have significantly advanced the state of the art in domains such as natural language processing and computer vision. These models can learn from large unannotated datasets and generalize across a wide range of downstream tasks. Notable early examples include BERT (Devlin et al., 2019) and GPT-3 (Brown et al., 2020) in natural language processing, as well as ViT (Dosovitskiy et al., 2020) and CLIP (Radford et al., 2021) in computer vision. These models often use self-supervised or unsupervised pretraining to learn rich, transferable representations. This approach has changed how models are built and used, removing the need for task-specific models and reducing the dependence on large labeled datasets for every single task. Inspired by these successes, there is growing interest now in extending the foundation models methodology to other modalities, including graphs.
@@
However, developing graph foundation models (GFMs) is much more challenging. Unlike text and images, graph data does not constitute a single domain. Instead, graphs are used to represent data from different domains, e.g., social networks (both virtual and real-world), information networks, transportation networks, co-purchasing networks, various physical, biological, or engineering systems, or even networks of abstract concepts. As a consequence, both the structure of a graph and its attributes (features and labels) may vary significantly across graph datasets and tasks. Moreover, the amount and diversity of the available graph data are significantly lower compared to those in computer vision and natural language processing. Taken together, these challenges call into question the very feasibility of graph foundation models.
@@
However, despite facing similar challenges, tabular machine learning has recently witnessed the emergence of the first successful tabular foundation models (TFMs) such as TabPFNv2 (Hollmann et al., 2025) or LimiX (Zhang et al., 2025a). Many of these models are based on the framework of prior-data fitted networks (PFNs, Müller et al., 2022; Hollmann et al., 2023). PFNs are pretrained on diverse synthetic datasets drawn from a prior to approximate Bayesian inference and can make predictions via in-context learning. Strong performance in the tabular domain suggests that PFNs offer a promising path towards building foundation models for both tabular and graph domains.
@@
The recent works G2T-FM (Eremeev et al., 2025), TAG (Hayler et al., 2025), and TabPFN-GN (Choi et al., 2025) have made the first step towards adopting PFNs for graph tasks by utilizing foundation models for tabular data to create graph foundation models for node-level tasks. For this, they augment node features with graph-based information such as neighborhood-aggregated features or Laplacian positional encodings. This allows for transforming a graph node-level prediction problem into a tabular prediction problem and applying an existing tabular foundation model to this task. The resulting approach shows strong performance, but such models still depend heavily on hand-crafted features and lack large-scale pretraining on diverse graph data. As a result, they are limited in their ability to capture complex graph patterns.
@@
In this work, we make the next step by proposing GraphPFN, a PFN-based model designed and pretrained specifically for graph node-level tasks. Following the PFN framework, we pretrain GraphPFN on synthetic datasets drawn from a carefully designed graph prior. For generating graph structures, we propose an approach that combines multiple stochastic block models and augments them with a preferential-attachment process. We then generate graph-structure-dependent node attributes for our graphs by augmenting tabular structured causal models (SCMs, Hollmann et al., 2023; Qu et al., 2025) typical for tabular PFNs with message-passing mechanisms at random SCM nodes. This method allows us to efficiently generate millions of realistic and diverse synthetic graph datasets. Then, we initialize GraphPFN from the tabular foundation model LimiX (Zhang et al., 2025a) and add an attention-based message-passing layer to each block of LimiX. This allows the model to learn complex graph-specific patterns while retaining its ability to handle diverse features and labels inherited from LimiX.
@@
Our experiments show that on diverse real-world graph node-level prediction datasets, GraphPFN achieves strong in-context learning performance, competitive with the best current models — well-tuned traditional GNNs (Kipf & Welling, 2017; Hamilton et al., 2017; Veličković et al., 2018; Shi et al., 2021) with improved architectures (Platonov et al., 2023b) and recent TFM-based GFMs G2T-FM (Eremeev et al., 2025) and TAG (Hayler et al., 2025). Furthermore, after finetuning, GraphPFN outperforms all other approaches, setting a new state of the art for the considered datasets.
@@
  • We propose GraphPFN, which is, to the best of our knowledge, the first publicly available PFN-based model designed and pretrained specifically for graph node-level tasks.
  • We introduce a novel graph prior for the efficient generation of realistic synthetic attributed graphs.
  • We demonstrate that the finetuned GraphPFN outperforms both strong traditional GNN baselines and existing graph foundation models.
@@ 2.2 Tabular foundation models
In their pioneering work TabPFN (Hollmann et al., 2023) and its successor TabPFNv2 (Hollmann et al., 2025), the authors proposed to utilize the framework of prior-data fitted networks to create tabular foundation models and showed that such models can achieve strong results, competitive with other approaches (Erickson et al., 2025). Nowadays, TFMs have become an active research area, with several new methods released recently (Zhang et al., 2025a; b; Grinsztajn et al., 2025; Qu et al., 2026). Some methods focus on scalability (Qu et al., 2025) or faster inference (Mueller et al., 2025), while others emphasize training on real-world datasets rather than synthetic tasks (Ma et al., 2025). Together, these works broaden the design space of tabular foundation models by trading off data sources, computational efficiency, and scalability.
@@ 2.3 Graph foundation models
Similar to foundation models for tabular data, graph foundation models face the challenge of handling datasets from diverse domains. A particularly difficult, yet crucial, aspect is managing the wide variety of node attributes (features and labels) present in different graphs. Early GFMs did not fully address this issue. They often relied on dimensionality reduction techniques such as principal component analysis or singular value decomposition (Xia & Huang, 2024; Zhao et al., 2024; Wang et al., 2025; Yu et al., 2025), or restricted their focus to graphs where node attributes are all textual (Wang et al., 2024; He et al., 2025; Liu et al., 2024).
More recent works, such as G2T-FM (Eremeev et al., 2025), TAG (Hayler et al., 2025), and TabPFN-GN (Choi et al., 2025), have explored leveraging tabular foundation models to better address feature diversity in graph datasets. These approaches incorporate hand-crafted features, for example, neighborhood-aggregated features or Laplacian positional encodings, to effectively convert graph information into tabular features. Empirical results show that these methods achieve strong results, always significantly outperforming prior GFMs (Xia et al., 2024; Xia & Huang, 2024; Zhao et al., 2024; Finkelshtein et al., 2025; Zhao et al., 2025) and frequently outperforming well-tuned GNNs trained from scratch (Kipf & Welling, 2017; Hamilton et al., 2017; Veličković et al., 2018; Shi et al., 2021) with improved architectures (Platonov et al., 2023b), supporting the utility of employing tabular foundation models as a basis for learning on graph data.
@@ 3.1 Architecture
Our model architecture extends the tabular foundation model LimiX (Zhang et al., 2025a) by adding attention-based message-passing layers to each of its transformer blocks. This strategy is inspired by recent findings that tabular foundation models can already learn patterns relevant for various graph tasks (Eremeev et al., 2025; Hayler et al., 2025; Choi et al., 2025). By initializing our model with a pretrained tabular foundation model instead of training from scratch, we leverage these learned representations, which significantly reduces computational costs while still achieving strong results.
Below, we first summarize the common architecture of a tabular foundation model (Hollmann et al., 2025; Zhang et al., 2025a), considering LimiX as the specific example, to clarify how GraphPFN represents samples (nodes) and features, and how attention flows in the base model. We then describe how the graph adapters modify this flow to leverage the graph topology.
@@
LimiX (Zhang et al., 2025a) is a transformer-style foundation model for tabular data that departs from the common design of representing each sample with a single fixed-length embedding. Instead, it uses a multi-token representation: for every sample, each feature contributes one token, [^1] In the current implementation, two features are grouped into one token, but we omit this detail for clarity. yielding a token grid with one axis for features and one for samples. This design naturally handles a variable number of heterogeneous features in different datasets without changing and retraining the model.
A LimiX transformer block contains three attention layers, each followed by an element-wise feed-forward network (FFN). Two layers are feature-level multi-head attention (MHA) modules that operate within a sample, allowing all feature tokens of the same sample to attend to each other. The third is a sample-level MHA that operates within a feature across samples, allowing tokens corresponding to the same feature to exchange information across the dataset. The feature-level MHAs enable rich interactions among features within each sample, while the sample-level MHA supports in-context learning by transferring information across samples for the same feature.
@@
To inject graph structure information without disrupting LimiX’s tokenization, we add a message-passing adapter that implements scaled dot product attention between neighboring nodes at the end of every LimiX transformer block (see Figure 1 for an illustration). Note that we use the scaled dot product attention that is identical to the original Transformer attention (Vaswani et al., 2017) except for being restricted to 1-hop graph neighborhoods. Intuitively, our message-passing adapter performs a second, graph-structure-aware round of sample-level attention: tokens may exchange information only along the observed edges. In contrast to the global sample-level attention of LimiX (which is masked by the PFN protocol), the graph adapter is masked by the adjacency and therefore routes information locally, from each node to its neighbors. Because we keep the per-feature token representation intact, the adapter runs over the sample axis for each feature token independently, using the same graph mask across features.
@@
Each adapter is implemented as a sparse, multi-head attention module with the adjacency as its mask, where two nodes attend to each other if and only if an edge connects them. Similar to other attention modules, it is followed by a feed-forward network independently applied to each token. Both the FFN and the message-passing layer are wrapped with residual connections (He et al., 2016) and layer normalization (Ba et al., 2016), mirroring the structure of the LimiX blocks for stable optimization.
@@ 3.2 Pretraining
GraphPFN was pretrained under the PFN framework (Müller et al., 2022; Hollmann et al., 2023; 2025) by continuing training from the LimiX checkpoint. In total, we used 2,240,0002{,}240{,}000 synthetic datasets generated according to the prior described below. Pretraining ran for approximately 7 days on 8 NVIDIA A100 80GB GPUs. Each GPU processed one synthetic dataset per step with 20 gradient accumulation steps, resulting in 160 datasets per optimizer step.
Inspired by TabICL (Qu et al., 2025), we pretrained GraphPFN in two stages to optimize training efficiency by gradually increasing the size of synthetic datasets. Specifically, Stage 1 ran for 10,00010{,}000 optimizer steps with sample sizes drawn log-uniformly from 1,0001{,}000 to 2,0002{,}000 and took approximately 2 days. Stage 2 ran for 4,0004{,}000 optimizer steps with sample sizes drawn log-uniformly from 8,0008{,}000 to 10,00010{,}000 and took approximately 5 days.
We optimized the model using AdamW (Loshchilov & Hutter, 2019) with weight decay $0.1$. We used a cosine annealing learning rate schedule (Loshchilov & Hutter, 2017) with linear warmup over the first $10\%$ of training steps and a base learning rate of $3\cdot 10^{-4}$. To improve training stability and preserve the feature modeling capabilities of LimiX, we froze all model layers except the graph adapters. Only these components were updated during pretraining. To further stabilize training, we applied an exponential moving average (EMA) to the weights throughout pretraining, with decay $0.98$ in Stage 1 and $0.95$ in Stage 2.
@@
We optimize a joint objective that combines the PFN supervised loss with the masked graph modeling (MGM) loss (Li et al., 2023). For the supervised PFN term, we sample a random set of context nodes for each dataset, compute predictions for all other nodes, and minimize the supervised loss on them. We use the cross-entropy loss for classification tasks and the mean squared error loss for regression tasks.
To encourage more graph-aware representations, inspired by the success of GNN self-supervised pretraining, we add the masked graph modeling term from Li et al. (2023). Specifically, we randomly sample a fraction $p=0.1$ of edges as positive samples, remove these edges from the input graph, and uniformly sample an equal number of unconnected node pairs as negative examples. We then train the model with the cross-entropy loss to distinguish between positive and negative edges. For this, we apply an additional MLP head to the pointwise multiplication of the target embeddings from the last layer for the source and destination nodes of each sampled edge. The total loss is the sum of the supervised loss and the MGM loss, with a coefficient of $0.1$ applied to the MGM term.
Learning curves
To monitor the progress of pretraining, we evaluated the in-context learning performance of GraphPFN on several datasets from the GraphLand benchmark (Bazhenov et al., 2025) (see Section 5.1 for the evaluation procedure) every 100 steps during the first stage and every 20 steps during the second stage. We emphasize that these evaluations were used only for a post-hoc analysis of training progress: test set performance was not used for early stopping or any other form of model selection. Figure 2 shows that the first stage enables GraphPFN to quickly reach reasonable performance on most datasets. On datasets such as artnet-views, tolokers-2, and twitch-views, where the performance gap between the graph-agnostic LimiX and GraphPFN is largest, most of the gap is closed within the first 1-3K gradient steps. Nevertheless, the second stage consistently improves performance across all datasets and is essential for achieving strong final performance of GraphPFN.
Figure 2: In-context learning performance of intermediate checkpoints of GraphPFN. The $x$-axis represents the number of steps, and the $y$-axis represents the metric on the test set. We also add vertical line at 10K steps to separate two pretraining stages.